A wire is formed into a circle having a diameter of 10.0 cm and placed in a uniform magnetic field of 3.00 mT. The wire carries a current of 5.00 A. Find (a) the maximum torque on the wire and (b) the range of potential energies of the wire–field system for different orientations of the circle. 


1
Expert's answer
2021-02-08T18:42:06-0500

"a) \tau_{max}=MB=IAB=l\pi r^{2}=5\times \pi"

"=375\pi \times 10^{-7}NM"

b) "U_{max}=+MB, U_{min}=-MB"